How fast is something traveling if it goes the distance of 1000m in 75 seconds?
Answer by Alan3354(69443)

How fast is something traveling if it goes the distance of 1000m in 75 seconds?
---------
d = r*t
r = d/t = 1000/75
r = 13.33... m/sec